### v2

# Testing the Homogeneity of Proportions for Correlated Bilateral Data via the Clayton Copula

Abstract
Handling highly dependent data is crucial in clinical trials, particularly in fields related to ophthalmology. Incorrectly specifying the dependency structure can lead to biased inferences. Traditionally, models rely on three fixed dependence structures, which lack flexibility and interpretation. In this article, we propose a framework using a more general model — copulas — to better account for dependency. We assess the performance of three different test statistics within the Clayton copula setting to demonstrate the framework’s feasibility. Simulation results indicate that this method controls type I error rates and achieves reasonable power, providing a solid benchmark for future research and broader applications. Additionally, we present analyses of two real-world datasets as case studies.
Keywords Homogeneity test, Paired correlated data, Copulas, Risk difference, Hypothesis testing
1 Introduction
In randomized clinical trials (RCTs), researchers are typically interested in evaluating treatment effects across various groups. By comparing each treatment group to a control group, the treatment effect for each group can be assessed. Additionally, comparing differences among treatment groups allows for further inferences about new study or treatment designs. Therefore, understanding the differences between groups is especially important in RCTs. However, in ophthalmologic studies, samples are often not based solely on individual patients but rather on paired eye samples. Some characteristics may show in one or both eyes, which increases the complexity of the model. In traditional likelihood-based methods, different outcome combinations result in different likelihood contributions while similar considerations arise in studies involving other paired or grouped organs, such as ears, legs, or fingers. In the context that follows, we use ophthalmologic studies as an example to model and analyze bilateral eye data.
In bilateral RCT data, a frequently addressed issue is the dependency. Failure to account for the dependence will lead to biased statistical inference (Rosner,, 1982; Dallal,, 1988; Donner and Banting,, 1988). Naturally, we assume that the data from both eyes of the same sample exhibit a certain degree of dependence. For example, there are inherent differences in corneal thickness between individuals, which leads to noticeable distinctions in the design of myopic laser surgeries. However, the corneal thickness of both eyes within the same individual is usually very similar, so the surgical design for both eyes of an individual is often alike. Ignoring the similarity between both eyes when making inferences from RCT data can easily lead to an overestimation of statistical power. Therefore, appropriately considering dependence is very important.
In previous studies, Rosner, (1982) proposed a "constant R" model to model intraclass correlation. The properties of this model and various related hypothesis testing methods were further investigated. For example, Ma et al., (2015) developed statistical tests for the homogeneity of multiple treatment groups. When the sample size is limited, exact methods were proposed by Liu et al., 2017a . However, the "constant R" model performed poorly because the characteristic is almost certain to occur bilaterally with widely varying group-specific prevalence. Dallal, (1988) developed a model that further employed a more flexible conditional probability assumption with group differences. Li et al., (2020) proposed asymptotic and exact methods for testing the homogeneity of the conditional probabilities among multiple groups. On the other hand, Donner, (1989) made assumptions based on a fixed value of correlation and proposed an adjusted Pearson’s chi-square test to test the differences among proportions arising from eye-specific binary data. Base on Donner’s model, homogeneity tests for comparing multiple treatment groups using asymptotic and exact methods were investigated by Ma and Liu, (2017) and Liu et al., 2017b , respectively.
In recent years, both clinical trials and observational studies have produced a growing amount of paired data. Many researchers have realized that paired data differ from other data types. As a result, this field has gained considerable attention and importance. Over a relatively short period, numerous studies examining different assumptions have emerged based on the three fundamental approaches. Liang et al., (2024) studied homogeneity tests of the risk difference between two groups across different strata and proposed confidence interval construction of the risk difference when there is no stratification effect, and Tian and Ma, (2024) proposed three statistical tests for assessing stratification effects on the risk ratio of two groups under Dallal’s model. Zhang and Ma, (2023) studied stratification effect on the risk difference between two groups under Donner’s model. Homogeneity tests of odds ratio was also examined by Hua and Ma, (2024) under different strata. Zhang and Ma, (2024) investigated many-to-one simultaneous confidence intervals of risk ratios.
We recognize that the correlation assumptions in these models are based on linear relationships. However, binary outcomes in eye data may involve more complex and potentially nonlinear dependence structures concerning proportion rates. Based on this, a more general framework capable of handling both linear and nonlinear dependencies warrants consideration. Copula functions are one of the popular tools of addressing both linear and nonlinear dependencies. Copula is also powerful to offer strong interpretative ability. It is widely employed in various research areas within clinical trials and observational studies, for instance, Emura and Chen, (2016), Emura et al., (2017), and Huang et al., (2021) applied bivariate copula functions to survival analysis for modeling semi-competing risks data. Modeling discrete data has been explored by Panagiotelis et al., (2012), Koopman et al., (2018), and Huang and Emura, (2024). Binary outcome models, a special case of discrete margin models, have been investigated in regression applications (Radice et al.,, 2016; Mesfioui et al.,, 2023), but their use in clinical data remains limited. Some well-established examples of the corresponding tools such as goodness-of-fit tests (Wang,, 2003; Huang and Emura,, 2021) and tests for symmetry (Jaser and Min,, 2021; Lyu and Belalia,, 2023) are available for reference. The prevalence and versatility of copula models allow this problem to be simplified to how to utilize copulas for handling binary outcome RCT data.
In this context, a copula can be understood as the underlying dependence structure of the treatment success rate/incidence rate model for both eyes. For example, it can be a linear correlation, such as the traditional Pearson’s $\rho$ model. Alternatively, copulas can allow tail dependence where there is strong dependency at low treatment success rates, but the dependency decreases as the success rate increases. Copulas, as a collection of these potential patterns, offer numerous options. Furthermore, the outcome data, that is, the marginal binary data, appear as paired samples representing these success rate/incidence rate models, containing potential dependence information. Therefore, it is reasonable to use copula models to generalize this application.
Our contributions are as follows:
  1. 1.
    We propose a novel framework to handle dependent binary outcomes of eyes in ophthalmology trials. This framework integrates a wide range of linear and nonlinear dependence structures by incorporating copulas, thereby achieving excellent flexibility. Several classical approaches (Rosner,, 1982; Dallal,, 1988; Donner,, 1989) can be generalized by our framework.
  2. 2.
    The proposed framework standardizes the correlation measure to Kendall’s tau under Archimedean copulas, facilitating the comparison of correlations across models.
  3. 3.
    We provide an example of performing likelihood inference and hypothesis testing for the treatment success rate parameters under our copula model. This inference forms the fundamental basis for future research on model comparison and goodness-of-fit test.
We primarily discuss the copula model and the rationale for using copulas with binary outcomes in Section 2.1. Sections 2.2 and 2.3 provide technical details on likelihood inference and test statistics. In Section 3, we outline the simulation design used to evaluate the statistical power of various test statistics within our proposed model. Section 4 presents a data analysis illustrating the model. Finally, in Section 5, we discuss our findings and outline future work based on this framework.
2 Method
We consider an ophthalmology dataset comparing $g$ groups of individuals, each with paired eye information. Group $i$ consists of $m_{i}$ individuals, totaling $N=\sum_{i=1}^{g}m_{i}$ observations. For each individual $j$ in group $i$, the status of the $k$th eye is represented by $Z_{ijk}$, where $Z_{ijk}=1$ indicates a diseased eye and $Z_{ijk}=0$ denotes a healthy eye, with $k=1,2$. To summarize the data (see Table 1), let $m_{i\ell}=\sum_{j=1}^{m_{i}}\mathbb{I}(Z_{ij1}+Z_{ij2}=\ell)$ represent the number of individuals with exactly $\ell$ diseased eyes in group $i$, and let $S_{\ell}=\sum_{i=1}^{g}m_{i\ell}$ denote the total number of individuals with $\ell$ diseased eyes across all groups.
Table 1: Frequencies of the Number of Diseased Eyes
Number of Diseased Eyes Group 1 Group 2 Group $g$ Total
0 $m_{10}$ $m_{20}$ $m_{g0}$ $S_{0}$
1 $m_{11}$ $m_{21}$ $m_{g1}$ $S_{1}$
2 $m_{12}$ $m_{22}$ $m_{g2}$ $S_{2}$
Total $m_{1}$ $m_{2}$ $m_{g}$ NN
2.1 Copula Model
The term “copula” originates from Latin, meaning link, tie, or bond. A bivariate copula is a function that links two uniform marginal distributions such that
$\Pr(U\leq u,V\leq v)=C_{\theta}(u,v),$
where $U$ and $V$ are the uniform random variables on $[0,1]$, and $\theta$ is the copula parameter. The joint distribution function satisfies properties as the usual joint probability such as $C(u,1)=u$, $C(1,v)=v$, and $C(u,0)=C(0,v)=0$. Copulas are advantageous thanks to Sklar’s theorem Sklar, (1959), which allows for different marginal distributions, enabling for arbitrary random variables $X$ and $Y$ such that
$\Pr(X\leq x,Y\leq y)=C_{\theta}(F_{X}(x),F_{Y}(y)).$
Notably, $C_{\theta}$ is unique when the marginals are continuous. Further details can be found in Nelsen, (2006).
Let $\Pr(Z_{ijk}=1)=\pi_{ik}$ denote the underlying treatment success rate of the $k$th eye in group $i$. The paired eye data can be modeled as
$\Pr(Z_{ij1}\leq z_{ij1},Z_{ij2}\leq z_{ij2})=C_{\theta}\bigl(F_{i1}(z_{ij1}),F_{i2}(z_{ij2})\bigr),$
where $C_{\theta}$ is a copula function. The marginal distributions follow a Bernoulli CDF:
$F_{ik}(z_{ijk})=\begin{cases}1-\pi_{ik},&\text{if }0\leq z_{ijk}<1,\\ 1,&\text{if }z_{ijk}\geq 1.\end{cases}$
Using copula properties, we can simplify the model algebraically. In total, there are four joint probabilities:
  1. 1.
    $\Pr(Z_{ij1}=0,Z_{ij2}=0)=C_{\theta}\bigl(1-\pi_{i1},1-\pi_{i2}\bigr)\overset{*}{=}C_{\theta}(1-\pi_{i},1-\pi_{i})$,
  2. 2.
    $\Pr(Z_{ij1}=1,Z_{ij2}=0)=1-\pi_{i2}-C_{\theta}\bigl(1-\pi_{i1},1-\pi_{i2}\bigr)\overset{*}{=}1-\pi_{i}-C_{\theta}(1-\pi_{i},1-\pi_{i})$,
  3. 3.
    $\Pr(Z_{ij1}=0,Z_{ij2}=1)=1-\pi_{i1}-C_{\theta}\bigl(1-\pi_{i1},1-\pi_{i2}\bigr)\overset{*}{=}1-\pi_{i}-C_{\theta}(1-\pi_{i},1-\pi_{i})$,
  4. 4.
    $\Pr(Z_{ij1}=1,Z_{ij2}=1)=1-\bigl(1-\pi_{i1}\bigr)-\bigl(1-\pi_{i2}\bigr)+C_{\theta}\bigl(1-\pi_{i1},1-\pi_{i2}\bigr)\overset{*}{=}1-2\bigl(1-\pi_{i}\bigr)+C_{\theta}(1-\pi_{i},1-\pi_{i})$.
[…]
The main advantage of using copulas is their flexibility in modeling dependence structures. Various copula families can be applied based on data patterns. For example, the Clayton copula is appropriate for lower-tail dependence, while the Joe copula is well-suited for upper-tail dependence. In Figure 1, we generate paired Bernoulli data using the Clayton copula,
$C_{\theta}(u,v)=\bigl(u^{-\theta}+v^{-\theta}-1\bigr)^{-1/\theta},\theta>0,$ (1)
for strong ($\tau=0.8$) and moderate ($\tau=0.5$) dependency with a marginal treatment success rate of 0.4. The points in the figures represent marginal uniform variables that reflect the copula’s correlation structure. Since the marginals are binary outcomes, a bubble plot effectively illustrates the frequencies of each case, where bubble sizes indicate the counts. The Clayton copula pattern shows stronger correlations for both successes treated eyes compared to both failure treated eyes. Cases with one success and one failure treated eye are less frequent under stronger correlation, and all data points cluster more closely around the center line.
Figure 1: Example of the Clayton Copula. The data points are generated from the Clayton copula. The dashed lines ($F_{1}=0.6$ and $F_{2}=0.6$) indicates the success treatment rate is $\pi_{i}=0.4$. Any value over these lines is categorized as $z_{ijk}=1$. Points are classified to the four different cases. The frequencies four each cases are shown by the bubbles.
The second advantage is the straightforward interpretation of the dependent parameters. Models proposed by Donner, (1989), Rosner, (1982), and Dallal, (1988) each provide different interpretations of dependency. Copula models offers Kendall’s tau measure that is free from the marginal distributions. Specifically, for Archimedean copulas, Kendall’s $\tau$ can be easily obtained from the generator functions Genest and Rivest, (1993). In the Clayton copula, $\tau=\frac{\theta}{\theta+2}$. Kendall’s $\tau$ is a unified measure of concordant pairs, thus the dependence is comparable between models. With the help of visualizing copulas, it offers a deeper understanding of dependence by capturing the specific structure beyond a single numerical value. The Pearson correlation coefficient between the treatment rates of two eyes is also available such that $\rho_{i}=\frac{C_{\theta}(1-\pi_{i},1-\pi_{i})-(1-\pi_{i})^{2}}{\pi_{i}(1-\pi_{i})}$, though it depends on the marginal distributions.
Our proposed copula model also generalizes several classical models. For example, Rosner, (1982) assumed $\Pr(Z_{ij1}=1\mid Z_{ij2}=1)=R\pi_{i}.$ where $R\pi_{i}$ is the treatment success rate. We can show that this is equivalent to the model with
$C_{\theta}(1-\pi_{i},1-\pi_{i})=R\pi_{i}^{2}-2\pi_{i}+1.$
This property holds in the independence case when $R=1$, where $C_{\theta}(1-\pi_{i},1-\pi_{i})=(1-\pi_{i})^{2},$ which implies an independence copula. Donner and Banting, (1988)’s model considered a common correlation coefficient $\rho=\text{Corr}(Z_{ij1},Z_{ij2})$. In the same concept, this model is denoted by copula as
$C_{\theta}(1-\pi_{i},1-\pi_{i})=(1-\rho)\pi_{i}^{2}+(\rho-2)\pi_{i}+1.$
The same generalization is applied to Dallal, (1988)’s model $\Pr(Z_{ij1}=1\mid Z_{ij2}=1)=\gamma_{i}$ leads to a copula
$C_{\theta}(1-\pi_{i},1-\pi_{i})=(\gamma_{i}-2)\pi_{i}+1.$
These results also reveal the relationship between these three classical models.
2.2 Likelihood Inference
We are interested in testing $H_{0}:$ $\pi_{1}=\pi_{2}=...\pi_{g}=\pi_{0}$ versus $H_{a}:$ $\pi_{i}\neq\pi_{j}$ for some $i\neq j$, where $\pi_{0}$ is an unknown quantity. Let $\tilde{M}=(m_{10},\dots,m_{g0},m_{11},\dots,m_{g1},m_{12},\dots,m_{g2})$ represent the observed summarized data. Assuming equal disease rates $\pi_{i}$ within group $i$, the log-likelihood function is expressed as:
[…]
$\quad+\,m_{i1}\log\left(2(1-\pi_{i})-2C_{\theta}(1-\pi_{i},1-\pi_{i})\right)$
$\quad+\,m_{i2}\log\left(1-2(1-\pi_{i})+C_{\theta}(1-\pi_{i},1-\pi_{i})\right).\Bigg]$
We apply the Clayton copula for inference. The Clayton copula function is defined as Equation 1 with $\theta>0$, thus the log-likelihood under the alternative becomes:
[…]
$\quad+\,m_{i1}\log\left(2(1-\pi_{i})-2\left(2(1-\pi_{i})^{-\theta}-1\right)^{-\frac{1}{\theta}}\right)$
$\quad+\,m_{i2}\log\left(1-2(1-\pi_{i})+\left(2(1-\pi_{i})^{-\theta}-1\right)^{-\frac{1}{\theta}}\right).\Bigg]$
Note that $1-2\pi_{i}\leq C_{\theta}\leq 1-\pi_{i}$ since $\Pr(Z_{ij1}=0,Z_{ij2}=0)$, $\Pr(Z_{ij1}=0,Z_{ij2}=1)$, $\Pr(Z_{ij1}=1,Z_{ij2}=0)$, and $\Pr(Z_{ij1}=1,Z_{ij2}=1)$ fall within the interval [0, 1]. It is straightforward to verify that the Clayton copula is an increasing function of $\theta$ given fixed $\pi_{i}$. Therefore, the following inequality holds:
$\lim_{\theta\to 0}(2(1-\pi_{i})^{-\theta}-1)^{-\frac{1}{\theta}}\leq C_{\theta}\leq\lim_{\theta\to\infty}(2(1-\pi_{i})^{-\theta}-1)^{-\frac{1}{\theta}},$
which is equivalent to
$(1-\pi_{i})^{2}\leq C_{\theta}\leq 1-\pi_{i}.$
It is clear that any $\theta>0$ satisfies the condition that $1-2\pi_{i}\leq C_{\theta}\leq 1-\pi_{i}$.
Under $H_{0}$, the log likelihood function can be expressed as:
[…]
$\quad+\,m_{i1}\log\left(2(1-\pi_{0})-2\left(2(1-\pi_{0})^{-\theta}-1\right)^{-\frac{1}{\theta}}\right)$
$\quad+\,m_{i2}\log\left(1-2(1-\pi_{0})+\left(2(1-\pi_{0})^{-\theta}-1\right)^{-\frac{1}{\theta}}\right)\Bigg].$
To obtain maximum likelihood estimators (MLEs) of $\pi_{i}$ and $\theta$, the Newton-type algorithm provided in the R function nlm() can be used (Dennis Jr and Schnabel,, 1996; Schnabel et al.,, 1985) since there is no analytical form of solutions. The algorithm finds a local minimizer of a real valued function without any constrains. However, all parameters in the proposed model are not necessarily from the real line. Hence, two transformations, $\log\left(\frac{\pi_{i}}{1-\pi_{i}}\right)$ and $\log(\theta)$, are employed to avoid introducing additional constrains. The final minimizer can be restored via a simple inverse of these functions. Denote the MLEs of $\pi_{i}$ and $\theta$ as $\hat{\pi}_{i}$ and $\hat{\theta}$ under $H_{a}$, respectively. Under the null hypothesis, the MLEs of $\pi_{0}$ and $\theta$ are represented by $\hat{\pi}_{H_{0}}$ and $\hat{\theta}_{H_{0}}$, respectively.
2.3 Hypothesis testing
2.3.1 Likelihood Ratio Test ($T_{LR}$)
The likelihood ratio (LR) test statistic, denoted as $T_{LR}$, is given by:
$).T_{LR}=2(\ell(\hat{\pi}_{1},\dots,\hat{\pi}_{g},\hat{\theta}\mid\tilde{M})-\ell_{0}(\pi_{0},\hat{\theta}\mid\tilde{M}$)).
Under $H_{0}$, $T_{LR}$ follows asymptotically a chi-square distribution with g1g-1 degrees of freedom (Wilks,, 1938).
2.3.2 Score Test ($T_{S}$)
Let $U=(\frac{\partial\ell}{\partial\pi_{1}},\frac{\partial\ell}{\partial\pi_{2}},...,\;\frac{\partial\ell}{\partial\pi_{g}},\frac{\partial\ell}{\partial\theta})^{T}$ and $I$ be the Fisher information matrix. The score test statistic is defined by:
$T_{S}=U^{T}I^{-1}U\big|\pi_{1}=\pi_{2}=...\pi_{g}=\hat{\pi}_{H_{0}},\theta=\hat{\theta}_{H_{0}},$
where
$I=\begin{bmatrix}I_{11}&0&&&0&I_{1(g+1)}\\ 0&I_{22}&&&0&I_{2(g+1)}\\ &&&\ddots&&\\ 0&0&&&I_{gg}&I_{g(g+1)}\\ I_{(g+1)1}&I_{(g+1)2}&&&I_{(g+1)g}&I_{(g+1)(g+1)}\end{bmatrix}_{(g+1)\times(g+1).}$
The expression of $I_{ii}$, $I_{i(g+1)}$, and $I_{(g+1)(g+1)}$ (i=1,2,,gi=1,2,...,g) can be found in the Appendix. The score test statistic $T_{S}$ can be further simplified as:
$I=\frac{-d_{i}(-\sum_{j\neq i}^{g}\frac{d_{i}I^{2}_{(g+1)j}}{I_{ii}I_{jj}}+\sum_{j\neq i}^{g}\frac{I_{(g+1)i}d_{j}I_{(g+1)j}}{I_{ii}I_{jj}}+\frac{d_{i}I_{(g+1)(g+1)}}{I_{ii}})}{\sum_{i=1}^{g}\frac{I^{2}_{(g+1)1}}{I_{ii}}-I_{(g+1)(g+1)}},$
where
[…]
$\text{with }\omega(\pi_{i})\;=\;{2}/{\bigl(1-\pi_{i}\bigr)^{\theta}}\;-\;1.$
$T_{S}$ is asymptotically distributed as a chi-square distribution with g1g-1 degrees of freedom according to Rao, (1948).
2.3.3 Wald Test ($T_{W}$)
To construct the Wald-type test, we first define $\beta=(\pi_{1},\pi_{2},...,\pi_{g},\theta)^{T}$ and note that the null hypothesis $H_{0}$ is equivalent to
$C\beta=\begin{bmatrix}1&-1&&&&0\\ &1&-1&&&0\\ &&\ddots&\ddots&&\vdots\\ &&&1&-1&0\end{bmatrix}_{(g-1)\times(g+1)}\times\begin{bmatrix}\pi_{1}\\ \pi_{2}\\ \vdots\\ \pi_{g}\\ \theta\end{bmatrix}=\begin{bmatrix}0\\ 0\\ \vdots\\ 0\end{bmatrix}_{(g-1)\times 1}.$
Then the Wald test statistic $T_{W}$ can be written as:
$T_{W}=(C\beta)^{T}(CI^{-1}C^{T})^{-1}(C\beta)|\pi_{1}=\hat{\pi}_{1},\pi_{2}=\hat{\pi}_{2},...,\pi_{g}=\hat{\pi}_{g},\theta=\hat{\theta}$,
which follows asymptotically a $\mathcal{X}^{2}$ distribution with g1g-1 degrees of freedom under the alternative hypothesis $H_{a}$ (Wald,, 1943).
3 Simulation
The type I errors (TIEs) and statistical powers of the three tests (LR test, score test, and Wald test) under our proposed copula model are investigated in this section using various parameter settings. Given a balanced design with the sample size $m_{i}=30$, $55$, and $100$, a random sample is drawn from different combination of $\pi_{i}$ and $\theta$, where $\pi_{i}\in\{0.4,0.5,0.6,0.7\}$ and $\theta\in\{0,2,8\}$. The higher the value of $\theta$, the stronger the correlation between the two eyes. Note that $\theta=0$ gives independence between two eyes. The empirical type I errors are determined by a total of 10,000 simulations and are presented in Table LABEL:tab:TIE3 and Table LABEL:tab:TIE6 under the group number $g=3$ and 6, respectively. To cover a broader parameter space, we first randomly generate a combination of $\pi_{i}$ and $\theta$ that makes $\Pr(Z_{ij1}=0,Z_{ij2}=0),\;\Pr(Z_{ij1}=0,Z_{ij2}=0),\;\Pr(Z_{ij1}=0,Z_{ij2}=0)$, and $\Pr(Z_{ij1}=0,Z_{ij2}=0)$ fall within the interval (0.1, 1]. The rationale of adding this restriction is to avoid sparse tables that lead to irregular behaviors of MLEs and test statistics. After that, ten thousand simulations are conducted to evaluate the empirical TIE. The aforementioned process is repeated 1,000 times and the results are presented in Figure 2 to Figure 5.
Table 2: The Empirical Type I Errors (%) when $g=3$ under the Nominal Level $\alpha=5\%$
$m_{i}=30$ $m_{i}=55$ $m_{i}=100$
$\theta$ $\pi_{i}$ $\rho_{i}$ $T_{LR}$ $T_{S}$ $T_{W}$ $T_{LR}$ $T_{S}$ $T_{W}$ $T_{LR}$ $T_{S}$ $T_{W}$
0 0.4 0 4.760 4.110 4.215 5.220 4.670 4.677 5.400 4.980 5.047
0.5 0 4.880 4.140 4.182 4.800 4.390 4.466 5.000 4.580 4.635
0.6 0 5.030 4.320 4.295 5.030 4.690 4.783 5.130 4.750 4.787
0.7 0 4.940 4.110 4.136 5.000 4.430 4.503 5.070 4.520 4.589
2 0.4 0.452 5.980 5.440 5.840 4.920 4.670 4.900 5.630 5.560 5.590
0.5 0.512 5.580 5.090 5.360 4.900 4.750 4.800 4.880 4.740 4.850
0.6 0.562 5.420 5.080 5.220 5.260 5.080 5.150 4.680 4.620 4.700
0.7 0.605 5.480 5.110 5.140 5.300 5.120 5.130 5.000 4.930 4.900
8 0.4 0.795 5.771 5.351 5.561 5.150 4.890 4.980 4.900 4.850 4.880
0.5 0.834 5.094 4.783 4.833 5.190 5.130 5.150 5.040 4.900 4.970
0.6 0.862 5.367 5.097 5.067 5.370 5.310 5.300 5.020 4.920 4.940
0.7 0.881 5.481 5.350 5.229 5.031 4.800 4.800 5.250 5.020 5.010
Table 3: The Empirical Type I Errors (%) when $g=6$ under the Nominal Level $\alpha=5\%$
$m_{i}=30$ $m_{i}=55$ $m_{i}=100$
$\theta$ $\pi_{i}$ $\rho_{i}$ $T_{LR}$ $T_{S}$ $T_{W}$ $T_{LR}$ $T_{S}$ $T_{W}$ $T_{LR}$ $T_{S}$ $T_{W}$
0 0.4 0 4.910 4.310 4.396 4.910 4.540 4.651 5.060 4.870 4.854
0.5 0 4.960 4.550 4.545 4.640 4.510 4.543 5.050 4.660 4.699
0.6 0 5.100 4.380 4.445 4.490 4.420 4.430 4.820 4.540 4.611
0.7 0 5.070 4.220 4.364 4.770 4.540 4.522 5.340 5.090 5.097
2 0.4 0.452 5.360 4.720 5.110 5.300 4.990 5.040 5.130 4.820 5.020
0.5 0.512 5.140 4.500 4.920 5.340 4.970 5.220 5.480 5.320 5.370
0.6 0.562 5.870 5.300 5.620 5.400 5.080 5.220 5.450 5.240 5.350
0.7 0.605 5.950 5.280 5.550 5.290 5.010 5.070 4.840 4.760 4.790
8 0.4 0.795 5.570 4.840 5.140 5.320 4.980 5.030 4.970 4.790 4.920
0.5 0.834 5.710 5.200 5.340 5.200 4.810 4.850 5.040 4.950 4.920
0.6 0.862 5.160 4.650 4.700 5.400 5.120 5.120 4.910 4.760 4.730
0.7 0.881 5.011 4.420 4.430 5.200 4.940 4.940 5.100 4.970 4.970
Figure 2: Violin Plots of Empirical Type I Errors for $g=3$ and $m_{i}=30$
Figure 3: Violin Plots of Empirical Type I Errors for $g=3$ and $m_{i}=100$
Figure 4: Violin Plots of Empirical Type I Errors for $g=6$ and $m_{i}=30$
Figure 5: Violin Plots of Empirical Type I Errors for $g=6$ and $m_{i}=100$
Next, we assess the power performance of the proposed methods under numerous alternative hypotheses defined in Table 4. The empirical powers when $g=3$ and 6 can be found in Table LABEL:tab:power3 and Table LABEL:tab:power6, respectively.
Table 4: Parameter Configurations for Empirical Powers
Case Group number $\pi_{i}$
1 3 $\pi_{1}=0.4$, $\pi_{2}=0.4$, $\pi_{3}=0.5$
2 3 $\pi_{1}=0.4$, $\pi_{2}=0.4$, $\pi_{3}=0.53$
3 3 $\pi_{1}=0.5$, $\pi_{2}=0.5$, $\pi_{3}=0.67$
4 3 $\pi_{1}=0.6$, $\pi_{2}=0.6$, $\pi_{3}=0.8$
A 6 $\pi_{1}=0.4$, $\pi_{2}=0.4$, $\pi_{3}=0.45$, $\pi_{4}=0.45$, $\pi_{5}=0.5$, $\pi_{6}=0.5$
B 6 $\pi_{1}=0.4$, $\pi_{2}=0.4$, $\pi_{3}=0.45$, $\pi_{4}=0.45$, $\pi_{5}=0.53$, $\pi_{6}=0.53$
C 6 $\pi_{1}=0.5$, $\pi_{2}=0.5$, $\pi_{3}=0.6$, $\pi_{4}=0.6$, $\pi_{5}=0.67$, $\pi_{6}=0.67$
D 6 $\pi_{1}=0.6$, $\pi_{2}=0.6$, $\pi_{3}=0.7$, $\pi_{4}=0.7$, $\pi_{5}=0.8$, $\pi_{6}=0.8$
Table 5: The Empirical Powers (%) when $g=3$ under the Nominal Level $\alpha=5\%$
$\theta$ Case Maximum of $m_{i}=30$ $m_{i}=55$ $m_{i}=100$
$|\pi_{i}-\pi_{j}|$, ($i\neq j$) $T_{LR}$ $T_{S}$ $T_{W}$ $T_{LR}$ $T_{S}$ $T_{W}$ $T_{LR}$ $T_{S}$ $T_{W}$
0 1 0.10 18.030 17.040 17.136 30.410 29.770 29.948 53.310 53.110 53.125
2 0.13 28.120 27.370 27.526 49.650 49.140 49.307 77.290 77.110 77.211
3 0.17 46.390 44.780 45.108 75.330 74.790 74.931 95.230 95.090 95.076
4 0.20 69.180 67.250 67.533 92.960 92.450 92.473 99.770 99.740 99.737
2 1 0.10 15.710 15.090 15.330 24.680 24.270 24.610 41.340 41.340 41.370
2 0.13 22.800 22.150 22.700 37.970 37.770 37.800 62.560 62.730 62.590
3 0.17 36.570 35.720 35.850 59.840 59.330 59.270 86.520 86.430 86.260
4 0.20 53.755 52.025 51.875 80.830 80.010 79.900 97.510 97.410 97.340
8 1 0.10 14.181 13.391 13.561 20.740 20.590 20.680 34.960 35.000 34.900
2 0.13 20.082 19.592 19.702 32.460 32.390 32.450 54.170 54.060 54.100
3 0.17 29.947 28.676 28.766 51.550 51.130 51.130 78.330 78.120 78.140
4 0.20 45.598 43.827 43.707 71.960 70.640 70.650 93.980 93.660 93.660
Table 6: The Empirical Powers (%) when $g=6$ under the Nominal Level $\alpha=5\%$
$\theta$ Case Maximum of $m_{i}=30$ $m_{i}=55$ $m_{i}=100$
$|\pi_{i}-\pi_{j}|$, ($i\neq j$) $T_{LR}$ $T_{S}$ $T_{W}$ $T_{LR}$ $T_{S}$ $T_{W}$ $T_{LR}$ $T_{S}$ $T_{W}$
0 A 0.10 17.600 16.370 16.480 31.350 30.720 30.812 56.290 55.740 55.743
B 0.13 29.206 27.576 27.932 52.730 52.230 52.492 82.830 82.660 82.745
C 0.17 51.100 49.200 49.590 80.180 79.930 80.096 98.290 98.260 98.238
D 0.20 74.870 73.220 73.433 96.310 96.050 96.164 99.980 99.980 99.980
2 A 0.10 15.320 14.020 14.710 23.960 22.880 23.700 43.040 42.240 42.710
B 0.13 23.360 21.800 22.420 40.120 39.200 39.620 68.450 68.040 68.180
C 0.17 37.820 36.060 36.640 64.490 63.620 64.230 91.120 90.910 91.120
D 0.20 57.506 55.586 55.916 86.680 86.180 86.290 99.010 98.980 98.990
8 A 0.10 13.580 12.610 12.840 20.240 19.560 19.760 35.270 34.800 34.950
B 0.13 19.430 18.300 18.560 34.170 33.440 33.540 58.870 58.490 58.550
C 0.17 31.930 30.290 30.420 55.010 54.250 54.290 83.620 83.420 83.460
D 0.20 48.277 46.013 46.023 77.260 76.340 76.320 97.180 97.070 97.070
4 Data Analysis
Two real-world examples are utilized to demonstrate the application of our proposed methods. The first is a cross-sectional study conducted in the Varamin district in Iran to assess the prevalence and causes of blindness and visual impairment (VI) (Rajavi et al.,, 2011). A total of 3,000 persons from 60 clusters were included using a multistage cluster systematic random sampling. Among them, 2819 persons were able to provide visual acuity measurements. The distribution of the available persons by age group and blindness status is exhibited in Table 7. The maximum likelihood estimates using the proposed method of parameters can be found in Table 8. The null hypothesis of homogeneous $\pi_{i}$ is rejected at $\alpha=5\%$ since $T_{LR}=136.589$ (p-value<0.0001), $T_{S}=178.749$ (p-value<0.0001), and $T_{W}=174.248$ (p-value<0.0001). The copula parameter estimate gives Kendall’s tau around 0.6 indicating positive dependence in visual acuity measurements.
Table 7: Example 1: Distribution of the Available Persons by Age Group and Blindness
Blindness 50–54 yrs 55–59 yrs 60–64 yrs 65–69 yrs 70–74 yrs 75–79 yrs $\geq$ 80 yrs Total
None 873 541 469 257 242 127 104 2613
Unilateral 23 17 18 16 32 30 29 165
Bilateral 2 8 4 5 3 9 10 41
Total 898 566 491 278 277 166 143 2819
Table 8: Maximum Likelihood Estimates of Parameters under $H_{0}$ and $H_{a}$ (Example 1)
Hypothesis MLE 50–54 yrs 55–59 yrs 60–64 yrs 65–69 yrs 70–74 yrs 75–79 yrs $\geq$ 80 yrs
$\hat{\pi}_{i}$ 0.015 0.030 0.027 0.048 0.067 0.139 0.163
$H_{a}$ $\hat{\theta}$ 4.581
$\hat{\rho}_{i}$ 0.065 0.120 0.109 0.180 0.236 0.395 0.434
$\hat{\pi}_{H_{0}}$ 0.044
$H_{0}$ $\hat{\theta}_{H0}$ 9.740
$\hat{\rho}_{H0}$ 0.301
Another example is from an observational study investigating the treatment effect of Orthokeratology (Ortho-k) on myopia at the First Affiliated Hospital of Xiamen University in 2023 (Liang et al.,, 2024). Ortho-k is a myopia correction method and is non-surgical. Subjects are required to wear specialized contact lenses overnight and the improvement of vision is temporary. There are two different lens designs. One is called corneal refractive therapy (CRT) and other is called vision shaping treatment (VST). Male subject distribution by design is presented in Table 9. The maximum likelihood estimates of parameters can be found in Table 10. We fail to reject the null hypothesis that $\pi_{VST}=\pi_{CRT}$ at the significant level of 5% as $T_{LR}=0.034$ (p-value=0.8546), $T_{S}=0.034$ (p-value=0.8543), and $T_{W}=0.034$ (p-value=0.8539). The copula parameter estimate gives strong positive dependence in vision outcomes.
Table 9: Example 2: Distribution of Male Subjects by Lens Design
No. of eyes with vision improvement VST CRT Total
0 11 6 17
1 4 2 6
2 3 2 5
Total 18 10 28
Table 10: Maximum Likelihood Estimates of Parameters under $H_{0}$ and $H_{a}$ (Example 2)
Hypothesis MLE VST CRT
$\hat{\pi}_{i}$ 0.276 0.303
$H_{a}$ $\hat{\theta}$ 3.051
$\hat{\rho}_{i}$ 0.466 0.491
$\hat{\pi}_{H_{0}}$ 0.286
$H_{0}$ $\hat{\theta}_{H0}$ 3.050
$\hat{\rho}_{H0}$ 0.475
5 Discussion
In this article, we proposed a new framework for testing the difference of event rates between groups with correlated paired data based on the Clayton copula. We investigated three hypothesis testing procedures and made likelihood inference. The optimization procedure is implemented using the R language. Simulation results indicated that all proposed testing procedures maintain satisfactory Type-I error control and exhibit reasonable power. Among them, the score test provides the best trade-off between error control and power, regardless of the number of groups, sample size, or parameter configurations. On the other hand, the LR test has inflated Type-I errors when the sample size is small, which is acceptable because real studies rarely involve extremely small samples. As the sample size increases, the three test procedures began to perform more similarly.
Compared to previous studies based on classical methods, our work yields similar simulation results. Therefore, we conclude that using copulas to model dependencies in correlated paired data under this setting is feasible. The success of this framework lays a foundation for our future research. In upcoming studies, selecting an appropriate copula and exploring properties of binary margin copulas will become increasingly important. Furthermore, this approach offers a more general perspective on similar models.
In addition, for specific copula models, we can extend our inference to confidence intervals. For instance, Tang et al., (2011) investigated asymptotic confidence intervals for the difference in disease rates between two groups, i.e., $\pi_{2}-\pi_{1}$. Yang et al., (2021) explored many-to-one simultaneous confidence intervals for comparing multiple treatments with a reference group. Pei et al., (2012) studied confidence intervals of proportion differences in a two-arm randomized clinical trial. Moreover, stratified designs can also be examined. For example, Qiu et al., 2019a ; Qiu et al., 2019b investigated asymptotic tests and confidence intervals for disease rates using stratified designs. Shen and Ma, (2018); Shen et al., (2019) studied tests for the homogeneity of differences between two disease rates across strata and provided interval estimations of a common risk difference. Beyond bilateral-only studies, we can further generalize to bilateral-mixed-with-unilateral designs. Under this combined data framework, for example, Ma and Wang, (2021) investigated asymptotic test methods for homogeneity of disease rates among multiple groups based on the “constant R” model.
Acknowledgment
We would like to thank Dr. Guanjie Lyu for his review and suggestions.
Conflict of Interest
All authors declare that they have no conflicts of interest.
Appendix A Fisher Information Matrix
The diagonal elements of the Fisher information matrix can be expressed as:
$I_{ii}=E\Bigl(-\frac{\partial\ell^{2}}{\partial\pi_{i}^{2}}\Bigr)=\frac{A_{i}}{B_{i}},$
where
$A_{i}=$ $4m_{i}+{\left(8m_{i}+2m_{i}{{\left(1-\pi_{i}\right)}}^{2\theta}\right)}{{\left(2-{{\left(1-\pi_{i}\right)}}^{\theta}\right)}}^{2/\theta}+{\left(8m_{i}{{\left(1-\pi_{i}\right)}}^{\theta}-2m_{i}{\left({{\left(1-\pi_{i}\right)}}^{2\theta+1}+6\right)}\right)}{{\left(2-{{\left(1-\pi_{i}\right)}}^{\theta}\right)}}^{1/\theta}$
$-8m_{i}{{\left(2-{{\left(1-\pi_{i}\right)}}^{\theta}\right)}}^{2/\theta}{{\left(1-\pi_{i}\right)}}^{\theta}$
$B_{i}=$ ${\left({{\left(2-{{\left(1-\pi_{i}\right)}}^{\theta}\right)}}^{1/\theta}-1\right)}{{\left({{\left(1-\pi_{i}\right)}}^{\theta}-2\right)}}^{2}{\left(\pi_{i}-1\right)}{\left(\pi_{i}-2\pi_{i}{{\left(2-{{\left(1-\pi_{i}\right)}}^{\theta}\right)}}^{1/\theta}+{{\left(2-{{\left(1-\pi_{i}\right)}}^{\theta}\right)}}^{1/\theta}-1\right)},\;$
$(i=1,2,...,g)$
and
$I_{(g+1)(g+1)}=E\Bigl(-\frac{\partial\ell^{2}}{\partial\theta^{2}}\Bigr)=\frac{C_{i}}{D_{i}},$
where
$C_{i}=$ $-m_{i}{{\left(2\log\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right)+2\theta\log\left(1-\pi_{i}\right)-\log\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right){{\left(1-\pi_{i}\right)}}^{\theta}\right)}}^{2}$
$\times{\left(3\pi_{i}{{\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right)}}^{1/\theta}-{{\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right)}}^{1/\theta}-2{\pi_{i}}^{2}{{\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right)}}^{1/\theta}+1\right)}$
$D_{i}=$ $\theta^{4}{{\left({{\left(1-\pi_{i}\right)}}^{\theta}-2\right)}}^{2}\Biggl[3\pi_{i}{{\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right)}}^{1/\theta}-2{{\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right)}}^{1/\theta}-3\pi_{i}{{\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right)}}^{2/\theta}$
$+{{\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right)}}^{2/\theta}+2{\pi_{i}}^{2}{{\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right)}}^{2/\theta}+1\Biggr].$
Elements at the bottom and right margins are:
$I_{i(g+1)}=I_{(g+1)i}=E\Bigl(-\frac{\partial\ell^{2}}{\partial\pi_{i}\partial\theta}\Bigr)=\frac{E_{i}}{F_{i}},$
where
$E_{i}=$ $-2m_{i}{\left(2\log\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right)+2\theta\log\left(1-\pi_{i}\right)-\log\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right){{\left(1-\pi_{i}\right)}}^{\theta}\right)}$
$\times\Biggl[\pi_{i}{(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}})}^{1/\theta}-{(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}})}^{1/\theta}$
$+\pi_{i}{(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}})}^{1/\theta}{{\left(1-\pi_{i}\right)}}^{\theta}-{\pi_{i}}^{2}{(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}})}^{1/\theta}{{\left(1-\pi_{i}\right)}}^{\theta}+1\Biggr]$
$F_{i}=$ $\theta^{2}{{\left({{\left(1-\pi_{i}\right)}}^{\theta}-2\right)}}^{2}{\left(\pi_{i}-1\right)}\Biggl[3\pi_{i}{{\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right)}}^{1/\theta}-2{{\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right)}}^{1/\theta}-3\pi_{i}{{\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right)}}^{2/\theta}$
$+{{\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right)}}^{2/\theta}+2{\pi_{i}}^{2}{{\left(-\frac{{{\left(1-\pi_{i}\right)}}^{\theta}-2}{{{\left(1-\pi_{i}\right)}}^{\theta}}\right)}}^{2/\theta}+1\Biggr],(i=1,2,...,g).$